A 0.20M solution of a weak monoprotic acid is 3.0% ionized. What is the approximate pH of the solution? show step by step solution. type solution. I won't give a 5 star rating if you scribble it down and i can't read your writing.

a) 11.0 to 12.0

b) 2.0 to 3.0

c) 4.0 to 5.0  

d) 9.0 to 10.0
Answer #1

Let the acid be HA and it dissociates as:

% ionisation = x*100/0.20

3.0 = x*100/0.20

x = 0.006 M

use:

pH = -log [H+]

= -log (0.006)

= 2.22

Answer: b